y = 5 when x = 2 . Average rate of change of y w.r.t. x between x = 2 and x = 5 is 3 and between x = 5 and x = 8 is 4. What is the value of y when x = 8


1
Expert's answer
2021-06-21T11:39:20-0400
"y(8)=y(5)+k'(8-5)\\y(5)=y(2)+k(5-2)\\y(5)=5+3(5-2)=14"

"y(8)=14+4(8-5)=26"